A snooker ball X of mass 200g moving at 20ms-1 to hit another stationary ball T of mass 400g. If after the collision Y moves off at 60o to the initial direction of X while X moves off at 15o to it's initial direction with a velocity of 12ms-1,. Calculate the final velocity of T


1
Expert's answer
2021-03-30T06:42:22-0400

To be given in question


Befor collision

Ball X

m1=200gram =0.2kg

u1=20m/sec 

Ball T


m2=400gram=0.4kg

u2=0m/sec 

After collision

 v1=12m/sec

"\theta" "1=60°
;\theta2=15°"


To be asked in question

V2=?

We take x-axis components

Momentum conservation law


Before collision=after collision


m1u1+m2u2=m1v1cos"\theta"1 +m2v2cos"\theta"2

Put value

"=0.2\times20+ 0.4\times0=0.2\times12\times cos60°+0.4\times V( 2)\times cos15°"

=4+0=2.4"\times" cos60°"+V(2)\times0.4\times cos15°"

4=1.2+V2"\times0.4 \times cos15°"

2.8=0.386 "\times" V2


V2=7.25m/sec